# Causal Inference for a Hidden Treatment

### v1

@@ 6 Extention to other functionals
In this section, we revisit the examples of §2 and derive corresponding influence functions within the hidden treatment model. Specifically, we continue to assume that the analyst has access to a surrogate $A$ and a proxy $Z$ for the hidden treatment $A^{*}$, that satisfy Assumptions 2, 3, and 4 We can leverage the same technique as in §3-4 to both identify these functionals and derive their influence functions. The detailed derivations are deferred to the Supplementary Materials. The estimation process follows in an analogous manner as above which we omit here for brevity.
For all following examples, we define
$q_{1}(A,Z,X)$ $=\frac{A-E(A\mid A^{*}=0,X)}{E(A\mid A^{*}=1,X)-E(A\mid A^{*}=0,X)}\frac{Z-E(Z\mid A^{*}=0,X)}{E(Z\mid A^{*}=1,X)-E(Z\mid A^{*}=0,X)},$ (20)
$q_{0}(A,Z,X)$ $=\frac{A-E(A\mid A^{*}=1,X)}{E(A\mid A^{*}=0,X)-E(A\mid A^{*}=1,X)}\frac{Z-E(Z\mid A^{*}=1,X)}{E(Z\mid A^{*}=0,X)-E(Z\mid A^{*}=1,X)}.$ (21)
@@
ETT (continued). An observed data influence function is derived by replacing the unobserved indicator $\mathbb{I}(A^{*}=1)$ in (3) with $q_{1}(A,Z,X)$ in (20) and replacing $\mathbb{I}(A^{*}=0)$ with $q_{0}(A,Z,X)$ in(21), which we can write as
$\phi_{ett}$ $=\frac{1}{\textrm{pr}(A^{*}=1)}q_{1}(A,Z,X)\{Y-E(Y\mid A^{*}=1,X)\}$
$\quad-\frac{1}{\textrm{pr}(A^{*}=1)}\frac{\textrm{pr}(A^{*}=1\mid X)}{1-\textrm{pr}(A^{*}=1\mid X)}q_{0}(A,Z,X)\{Y-E(Y\mid A^{*}=0,X)\}$
@@
The influence function in the observed data model is obtained by substituting $\mathbb{I}(A^{*}=a^{*})$ in (4) with $q_{a^{*}}(A,Z,X)$ in (20) or (20), depending on the value of $a^{*}$, which gives
$\phi_{a^{*},\gamma}=q_{a^{*}}(A,Z,X)\frac{\mathbb{I}(Y\leq\theta_{1}^{a^{*}})-\eta_{1}(X;\theta_{1}^{a^{*}})}{\eta_{2}(X)}+\eta_{1}(X;\theta_{1}^{a^{*}})-\gamma.$
@@
Semi-linear regression model (continued). The influence function in (5) can be reformulated as
@@
The influence function in the observed data model can be derived by replacing $\mathbb{I}(A^{*}=1)$ with $q_{1}(A,Z,X)$ defined in (20) and replacing $\mathbb{I}(A^{*}=0)$ with $q_{0}(A,Z,X)$ defined in (21), which gives
$\phi_{sl}$ $=q_{1}(A,Z,X)h(X)\{Y-\beta_{a^{*}}-\beta_{x}(X)\}\{1-\textrm{pr}(A^{*}=1\mid X)\}$
$\quad+q_{0}(A,Z,X)h(X)\{Y-\beta_{x}(X)\}\{-\textrm{pr}(A^{*}=1\mid X)\}.$
@@
Marginal structural generalized linear model (continued). The influence function in (6) can be rewritten as
@@
The influence function in the observed data model can be derived by replacing $\mathbb{I}(A^{*}=1)$ with $q_{1}(A,Z,X)$ defined in (20) and replacing $\mathbb{I}(A^{*}=0)$ with $q_{0}(A,Z,X)$ defined in (21), which gives
$\phi_{msm}$ $=\frac{q_{1}(A,Z,X)}{\textrm{pr}(A^{*}=1\mid X)}h(1,V)\left\{Y-E(Y\mid A^{*}=1,X)\right\}$
$\quad+\frac{q_{0}(A,Z,X)}{\textrm{pr}(A^{*}=0\mid X)}h(0,V)\left\{Y-E(Y\mid A^{*}=0,X)\right\}$
@@ 8 Real data example
Alzheimer’s Disease (AD) is a progressive neurodegenerative disorder that primarily affects the elderly, characterized by significant cognitive decline and memory impairment. One of the hallmark changes observed in AD is the atrophy of the hippocampus, a critical brain region involved in memory formation and spatial navigation. As the disease advances, neuronal damage occurs, resulting in brain atrophy that is especially pronounced in the hippocampus during the early stages of AD (Fox et al., 1996; Yu et al., 2022). We aim to analyze the impact of AD on hippocampal volume using data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI). The data usage acknowledgment is included in the Supplementary Materials.
@@
Our results indicate that for females, AD leads to a decrease of 1,756 mm3 in hippocampal volume, with 95% confidence interval of $[1,370,2,142]$. For males, AD leads to a decrease of 1,406 mm3 in hippocampal volume, with 95% confidence interval of $[853,1,960]$ (Fox et al., 1996). These findings appear to be aligned with prior research indicating that females are affected by AD more severely than males, and AD-control brain weight differences were significantly greater for females (Filon et al., 2016). For comparison, we also performed linear regression for both groups, with hippocampal volume as the outcome, error-prone diagnosis as the main exposure, age, and education years as regressors. These analyses indicate an AD-related decrease of 1,162 mm3 for females (95% CI: $[976,1,347]$) and 971 mm3 for males (95% CI: $[793,1,151]$), which are of smaller magnitude than our estimates, suggesting the potential for effect attenuation, a common consequence of exposure measurement error.
@@ 9 Discussion
In this paper, we considered the fundamental problem of causal inference with a hidden binary treatment, an all too common problem in epidemiology and the social sciences. We proposed a unified solution which delivers nonparametric identification of a rich class of causal parameters, by drawing on a strategy which carefully leverages a surrogate and a proxy variable of the hidden treatment to restore identification, while bypassing the need for validation data or restrictive functional form assumptions prominent in prior measurement error literature. An important contribution of the paper is to develop formal semiparametric theory which gives rise to influence function-based nonparametric estimators with important robustness and efficiency properties. The approach provides a generic roadmap for constructing novel semiparametric estimators in a large class of hidden treatment models. The paper also addresses the challenging issue of estimating high-dimensional nuisance functions which are defined in terms of the hidden treatment, therefore ruling out direct application of standard machine learning and nonparametric estimators. To resolve this difficulty, we proposed a semiparametric EM algorithm for estimating nuisance functions. We establish the semiparametric efficiency bound for binary outcomes, leaving the more general case for future research.

### v2

@@ 6 Extention to other functionals
In this section, we revisit the examples of §2 and derive corresponding influence functions in the hidden treatment model. Specifically, we continue to assume that the analyst has access to a surrogate $A$ and a proxy $Z$ for the hidden treatment $A^{*}$, that satisfy Assumptions 2, 3, and 4. We establish that one can readily leverage the same techniques as in §3-4 for identification, and we generically denote as $\psi^{*}$ the corresponding observed data functional for which we derive influence functions. The estimation of nuisance functions then follows in an analogous manner as above, details of which we omit for brevity. To present our results in a unified fashion, we note that all examples given in §2 have full data influence functions of the bi-linear form
$\begin{split}\phi_{\mathrm{g}}^{F}(Y,A^{*},X;\psi^{*})&=\sum_{a^{*}=0,1}IF_{a^{*};\psi^{*}}=\sum_{a^{*}=0,1}\Bigl{\{}q_{0,a^{*}}(X)h_{0,a^{*}}(X)g_{1,a^{*}}(Y,X)\mathbb{I}(A^{*}=a^{*})\\ &\quad+q_{0,a^{*}}(X)g_{2,a^{*}}(Y,X)\mathbb{I}(A^{*}=a^{*})+h_{0,a^{*}}(X)g_{3,a^{*}}(Y,X)+g_{4,a^{*}}(Y,X)\Bigr{\}},\end{split}$ (21)
where $q_{0,a^{*}}(X)$ and $h_{0,a^{*}}(X)$ are unknown nuisance functions, $g_{j,a^{*}}$ are known functions of $(Y,X)$, and $IF_{a^{*};\psi^{*}},a^{*}=0,1$, belongs to the doubly robust class of influence functions studied in Robins et al. (2008). To simplify the exposition, we suppress the influence function dependence on $\psi^{*}$ throughout as $E[IF_{a^{*};\psi^{*}}]=0,a^{*}=0,1$. The next theorem characterizes for each influence function in the class above, a corresponding observed data influence function in the hidden treatment model. While we state the result for the class given above, the proof is more general as it considers an even larger class of doubly robust influence functions introduced in Ghassami et al. (2022), an important generalization which includes influence functions for causal functionals identified using the proximal causal inference framework (Miao et al., 2016, 2018; Tchetgen Tchetgen et al., 2024) that do not strictly fall within the above doubly robust class (Cui et al., 2023). The proof is deferred to the Supplementary Materials §LABEL:supp-sec:proof_extension.
Theorem 6.
Suppose that the influence functions for the parameter of interest $\psi$ in the full data model $\mathcal{M}^{*}$ are of the form in (21), indexed by functions $\mathrm{g}=(g_{1,a^{*}},g_{2,a^{*}},g_{3,a^{*}},g_{4,a^{*}},a^{*}=0,1)\in\mathcal{G}$ for a certain space of functions $\mathcal{G}$ in $L_{0}^{2}(X,Y)$ implied by the semiparametric model; then
$\begin{split}\phi_{\mathrm{g}}&(Y,A,Z,X;\psi^{*})=\sum_{a^{*}=0,1}\Bigl{\{}q_{0,a^{*}}(X)h_{0,a^{*}}(X)g_{1,a^{*}}\left(Y,X\right)M_{a^{*},A}M_{a^{*},Z}\\ &+q_{0,a^{*}}(X)g_{2,a^{*}}\left(Y,X\right)M_{a^{*},A}M_{a^{*},Z}+h_{0,a^{*}}(X)g_{3,a^{*}}(Y,X)+g_{4,a^{*}}(Y,X)\Bigr{\}}\end{split}$ (22)
is an influence function for $\psi$ in the hidden treatment model $\mathcal{M}$, where
$\begin{split}M_{a^{*},A}=\frac{A-E(A\mid A^{*}=1-a^{*},X)}{E(A\mid A^{*}=a^{*},X)-E(A\mid A^{*}=1-a^{*},X)},\\ M_{a^{*},Z}=\frac{Z-E(Z\mid A^{*}=1-a^{*},X)}{E(Z\mid A^{*}=a^{*},X)-E(Z\mid A^{*}=1-a^{*},X)}.\end{split}$ (23)
The set of influence functions in $\mathcal{M}$ is $\{\phi_{\mathrm{g}}:\mathrm{g}\in\mathcal{G}\}+\Lambda^{\bot}$, where $\Lambda^{\bot}$ is given in (14).
Upon deriving the set of influence functions, an efficient estimator for $\psi^{*}$ in the hidden treatment model can be constructed as follows:
  1. 1.
    Obtain a preliminary (potentially inefficient) estimator $\hat{\psi}$ that solves $0=\mathbb{P}_{n}\hat{\phi}_{\mathrm{g}}(Y,A,Z,X)$ by using an estimated influence function $\hat{\phi}_{\mathrm{g}}(Y,A,Z,X)$ as an estimating function for a fixed choice $\mathrm{g}$.
  2. 2.
    Perform a one-step update in the direction of estimated efficient influence function
    $\hat{\psi}^{eff}=\hat{\psi}+\mathbb{P}_{n}\hat{\phi}_{eff}(Y,A,Z,X)$
    where $\hat{\phi}_{eff}(Y,A,Z,X)$ is an estimator of the efficient influence function of ${\phi}_{eff}(Y,A,Z,X)=\Pi(\phi_{\mathrm{g}}\mid\Lambda)$ evaluated at $\hat{\psi}$ given in (LABEL:supp-eq:est_eif) in the Supplementary Materials §LABEL:supp-sec:closed_form_w.
@@
ETT (continued). An observed data influence function is derived by replacing the unobserved indicator $\mathbb{I}(A^{*}=a^{*}),a^{*}=0,1$ in (3) with $M_{a^{*},A}(A,X)M_{a^{*},Z}(Z,X)$ in (23), which we can write as
$\phi_{ett}$ $=\frac{1}{\textrm{pr}(A^{*}=1)}M_{1,A}(A,X)M_{1,Z}(Z,X)\{Y-E(Y\mid A^{*}=1,X)\}$
$\quad-\frac{1}{\textrm{pr}(A^{*}=1)}\frac{\textrm{pr}(A^{*}=1\mid X)}{1-\textrm{pr}(A^{*}=1\mid X)}M_{0,A}(A,X)M_{0,Z}(Z,X)\{Y-E(Y\mid A^{*}=0,X)\}$
@@
The influence function in the observed data model is obtained by substituting $\mathbb{I}(A^{*}=a^{*})$ in (4) with $M_{a^{*},A}(A,X)M_{a^{*},Z}(Z,X)$ in (23), which gives
$\phi_{a^{*},\gamma}=M_{a^{*},A}(A,X)M_{a^{*},Z}(Z,X)\frac{\mathbb{I}(Y\leq\theta_{1}^{a^{*}})-\eta_{1}(X;\theta_{1}^{a^{*}})}{\eta_{2}(X)}+\eta_{1}(X;\theta_{1}^{a^{*}})-\gamma.$
@@
Semi-linear structural regression model (continued). The full data influence function in (5) can be reformulated as
@@
An influence function in the observed data model can be obtained by replacing $\mathbb{I}(A^{*}=a^{*}),a^{*}=0,1$ with $M_{a^{*},A}(A,X)M_{a^{*},Z}(Z,X)$ in (23), which gives
$\phi_{sl}$ $=M_{1,A}(A,X)M_{1,Z}(Z,X)h(X)\{Y-\beta_{a^{*}}-\beta_{x}(X)\}\{1-\textrm{pr}(A^{*}=1\mid X)\}$
$\quad+M_{0,A}(A,X)M_{0,Z}(Z,X)h(X)\{Y-\beta_{x}(X)\}\{-\textrm{pr}(A^{*}=1\mid X)\}.$
@@
Marginal structural generalized linear model (continued). The full data influence function in (6) can be rewritten as
@@
An influence function in the observed data model can be derived by replacing $\mathbb{I}(A^{*}=a^{*}),a^{*}=0,1$ with $M_{a^{*},A}(A,X)M_{a^{*},Z}(Z,X)$ in (23), which gives
$\phi_{msm}$ $=\frac{M_{1,A}(A,X)M_{1,Z}(Z,X)}{\textrm{pr}(A^{*}=1\mid X)}h(1,V)\left\{Y-E(Y\mid A^{*}=1,X)\right\}$
$\quad+\frac{M_{0,A}(A,X)M_{0,Z}(Z,X)}{\textrm{pr}(A^{*}=0\mid X)}h(0,V)\left\{Y-E(Y\mid A^{*}=0,X)\right\}$
@@ 8 Real data example
Alzheimer’s Disease (AD) is a progressive neurodegenerative disorder that primarily affects the elderly, characterized by significant cognitive decline and memory impairment. One of the hallmark changes observed in AD is the atrophy of the hippocampus, a critical brain region involved in memory formation and spatial navigation. As the disease advances, neuronal damage occurs, resulting in brain atrophy that is especially pronounced in the hippocampus during the early stages of AD (Fox et al., 1996; Yu et al., 2022). We aim to analyze the impact of AD on hippocampal volume using data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI). The data usage acknowledgment is included in §LABEL:supp-sec:adni_acknowledgment of the Supplementary Materials.
@@
Our results indicate that for females, AD leads to a decrease of 1,756 mm3 in hippocampal volume, with 95% confidence interval of $[1,370,2,142]$. For males, AD leads to a decrease of 1,406 mm3 in hippocampal volume, with 95% confidence interval of $[853,1,960]$ (Fox et al., 1996). These findings appear to be aligned with prior research indicating that females are affected by AD more severely than males, and AD-control brain weight differences were significantly greater for females (Filon et al., 2016). For comparison, we also performed linear regression for both groups, with hippocampal volume as the outcome, error-prone diagnosis as the main exposure, age, and education years as regressors. These analyses indicate an AD-related decrease of 1,162 mm3 for females (95% CI: $[976,1,347]$) and 971 mm3 for males (95% CI: $[793,1,151]$), which are of smaller magnitude than our estimates, suggesting the potential for effect attenuation, a common consequence of exposure measurement error. In addition to linear regression, we also implemented the more complex doubly robust Targeted Maximum Likelihood Estimator (TMLE) (Benkeser et al., 2017), which is essentially biased in this context and yields results similar to those of linear regression.
@@ 9 Discussion
In this paper, we considered the fundamental problem of causal inference with a hidden binary treatment, an all too common problem in epidemiology and the social sciences. We proposed a unified solution which delivers nonparametric identification of a rich class of causal parameters, by drawing on a strategy which carefully leverages a surrogate and a proxy variable of the hidden treatment to restore identification, while bypassing the need for validation data or restrictive functional form assumptions prominent in prior measurement error literature. An important contribution of the paper is to develop formal semiparametric theory which gives rise to influence function-based nonparametric estimators with important robustness and efficiency properties. The approach provides a generic roadmap for constructing novel semiparametric estimators in a large class of hidden treatment models. The paper also addresses the challenging issue of estimating high-dimensional nuisance functions which are defined in terms of the hidden treatment, therefore ruling out direct application of standard machine learning and nonparametric estimators. To resolve this difficulty, we proposed a semiparametric EM algorithm for estimating nuisance functions. We establish the semiparametric efficiency bound for discrete outcomes, and provide an approximation for continuous outcomes.